### v1

@@ 2.2 KNO scaling within QCD jets
In the asymptotic limit $Q\to\infty$, KNO scaling in multiplicity distributions within QCD jets naturally follows the definition of the GFs Konishi:1979cb ; Bassetto:1979nt ; Dokshitzer:1982ia ; Dokshitzer:1991wu ; Dremin:2000ep . As previously verified in ref. Duan:2025ngi , $\bar{n}_{a}(Q)P_{a}(n,Q)$ for partons within DLA, given in eq. (2.1.2), indeed converge to universal scaling functions for quark and gluon jets respectively for $Q\gg Q_{0}$. They are, however, quite different from those for hadrons in PYTHIA simulations and the LHC data. In this subsection, we evaluate the KNO scaling functions within MDLA for both quark and gluon jets, improving the DLA results Dokshitzer:1982ia ; Bassetto:1987fq ; Dokshitzer:1991wu by imposing energy conservation Dokshitzer:1993dc ; Dokshitzer:2025owq ; Dokshitzer:2025fky .
@@ 2.2.1 Definition of KNO scaling functions for quark and gluon jets
Following refs. Dokshitzer:1982ia ; Dokshitzer:1991wu , we first outline the definition of, and the method for evaluating, the KNO scaling functions for QCD jets, which holds exactly in the asymptotic limit $Q\to\infty$. Setting $u=\exp\{-\beta/\bar{n}_{a}\}$ in the definition of GFs in eq. (2) and taking the limit $Q\to\infty$ in $Z_{a}$ yields
@@
[…]
@@ 2.2.2 KNO scaling functions within MDLA
As we are interested in the large $Q$ limit, $n^{(m)}_{a}(Q)$ can be expressed in terms of $f_{a}^{(m)}$ and $\bar{n}_{a}(Q)$ according to eq. (28) in the above equation. Following ref. Dokshitzer:1993dc , we consider only the corrections arising from energy conservation, while neglecting the effects of the running coupling and higher-order corrections to the anomalous dimension $\gamma_{0}$. In this case, one can approximate the mean multiplicities as
@@
where the digamma function $\psi(z)\equiv\Gamma^{\prime}(z)/\Gamma(z)$. Following ref. Dokshitzer:1993dc , we neglect the $\psi$ terms. Then, starting with $f_{a}^{(0)}=f_{a}^{(1)}=1$, one can iteratively obtain, for $m\geq 2$,
$f^{(m)}_{g}$ $=\frac{\gamma_{0}m}{m^{2}-1}\sum\limits_{k=1}^{m-1}C_{m}^{k}\frac{\Gamma(\gamma_{0}k)\Gamma(\gamma_{0}(m-k)+1)}{\Gamma(\gamma_{0}m+1)}f^{(k)}_{g}f^{(q-k)}_{g}$,
@@
The DLA results Dokshitzer:1982ia ; Dokshitzer:1991wu are recovered in the limit $\gamma_{0}\to 0$.
Given the expressions $f_{a}^{(m)}$ above, the KNO scaling functions $\Psi_{a}(x)$ can be presented using an expansion in Laguerre polynomials Bassetto:1987fq , which obey the relations
@@
Figure 1: KNO scaling functions for quark and gluon jets. Left: Verification of KNO scaling within DLA by directly evaluating $P_{a}(n)$ according to eq. (2.1.2), using a fixed coupling $\alpha_{s}=0.2$, over a $p_{T}$ range from 0.1 to 2.5 TeV (solid), and comparing them to the DLA KNO scaling functions derived in the asymptotic limit $Q\to\infty$ (dashed). Within this $p_{T}$ range, we observe that the curves of $\bar{n}_{a}P_{a}(n)$ for different $p_{T}$ bins still show some dependence on the value of $\alpha_{s}$. With the chosen value, they all converge to the DLA KNO scaling functions. Right: Comparison of the MDLA KNO scaling functions evaluated from eq. (30) (solid) with the QCD-inspired forms in ref. Dokshitzer:2025fky (dashed) for $\gamma_{0}=0.4=\gamma$ and $c_{q}=2/3$.
@@
The left panel of figure 1 verifies the validity of KNO scaling in DLA with a fixed $\alpha_{s}$, which is assumed in the preceding MDLA derivation. This is verified by directly evaluating $P_{a}(n)$, as given in eq. (2.1.2), over a broad $p_{T}$ range from 0.1 to 2.5 TeV. With running coupling, the resulting curves of $\bar{n}_{a}P_{a}(n)$ all converge to some universal functions for quark and gluon jets, respectively, though with slight deviations from the asymptotic DLA scaling functions Duan:2025ngi . With fixed coupling, we find that $\bar{n}_{a}P_{a}(n)$ shows some dependence on the value of $\alpha_{s}$ and the results with $\alpha_{s}=0.2$ converge to the DLA scaling functions, as shown in this plot. Here, the asymptotic KNO functions within DLA are given by eq. (36) with $f_{a}^{(n)}$ evaluated in the limit $\gamma_{0}\to 0$, which agree with the results of refs. (Bassetto:1987fq, ; Dokshitzer:1993dc, ).
@@
The right panel of figure 1 compares the MDLA scaling functions evaluated above with the QCD-inspired expressions recently proposed in ref. Dokshitzer:2025fky . The comparison is shown for $\gamma_{0}=0.4$ Dokshitzer:1993dc (with $\gamma=0.4$ in the results of ref.Dokshitzer:2025fky ) and $c_{q}=2/3$ Dokshitzer:2025fky . For both quark and gluon jets, including energy conservation significantly shifts the peak of the scaling functions toward larger values of $x$. These significant discrepancies highlight the need to go beyond DLA to obtain reliable KNO scaling functions. Our results, obtained using Laguerre polynomial expansions, agree with the expression for gluon jets given in ref. Dokshitzer:1993dc , except in the $x\ll 1$ region, where our method may yield negative values, like the DLA results in ref. Bassetto:1987fq . Hence, for the remainder of the paper, we adopt the analytic form from ref. Dokshitzer:1993dc for gluon jets. In comparison with recently suggested forms in ref. Dokshitzer:2025fky , our results for gluon jets are slightly different from that at $x\lesssim 1$, starting around the peak region, while the differences for quark jets are more pronounced. We find that this difference can be traced back to the fact that the coefficients $f_{q}^{(m)}$, calculated iteratively above, no longer satisfy the relation within DLA: $\Phi_{q}(\beta)=[\Phi_{g}(\beta/c_{q})]^{c_{q}}$, which is used to obtain the quark-jet KNO function in refs. Dokshitzer:2025owq ; Dokshitzer:2025fky .
@@ 2.3 Comparison with KNO scaling functions using PYTHIA
As shown by the results above, while the mean parton multiplicity and the parton multiplicity distribution functions exhibit a logarithmic dependence on $Q_{0}$, the asymptotic KNO scaling functions remain infrared safe. Consequently, one may not expect logarithmically enhanced corrections when transitioning from the parton level to the hadron level. In this subsection, we compare the above results at the parton level with those for charged particles extracted in PYTHIA simulations Bierlich:2022pfr , which exhibit KNO scaling to a very reasonable level in both quark and gluon jets Duan:2025ngi .
@@
In our PYTHIA simulations of $pp$ collisions Duan:2025ngi , the event selection procedure follows the methodology in ref. ATLAS:2019rqw . Particles satisfying $p_{T}>0.5$ GeV and $|\eta|<2.5$ are used to reconstruct the two leading jets via the anti-$k_{t}$ algorithm Cacciari:2008gp with a radius parameter $R=0.4$, implemented in the FastJet package Cacciari:2011ma . The jets are required to satisfy $|\eta|<2.1$, and the leading and subleading jets must have a transverse momentum ratio $p_{T}^{\text{lead}}/p_{T}^{\text{sublead}}<1.5$. The PYTHIA samples are generated using the A14 tune, which has been shown to provide a good description for charged-particle multiplicity distribution observed by ATLAS ATLAS:2019rqw . In our analysis, the charged particles from jet constituents are used to define the multiplicity distributions, while jet $p_{T}$ is reconstructed from all particles inside the jet. In addition, quark and gluon jets are identified based on their spatial proximity to the initiating parton from the underlying $2\rightarrow 2$ hard scattering event.
@@
Figure 2: KNO scaling functions for quark and gluon jets in PYTHIA and MDLA. Left: Charged-particle multiplicity distributions for quark and gluon jets (solid) versus MDLA results (dashed) with $\gamma_{0}=0.43$ and $c_{q}=0.8$. Right: Negative binomial distributions (NBD, solid), with $k=5.4$ for quark jets and $k=8.5$ for gluon jets (see eq. (38)) compared to the same MDLA results (dashed).
@@
The left panel of figure 2 compares our MDLA results with PYTHIA simulations that include both MPI and hadronization, facilitating comparison with ATLAS data in the next sections. We find that incorporating energy conservation indeed improves significantly the agreement between the KNO scaling functions and the PYTHIA results for both quark and gluon jets. Inspired by refs. Dokshitzer:2025owq ; Dokshitzer:2025fky , we take $c_{q}$ as a fitting parameter rather than fixing it to the DLA value $C_{F}/C_{A}$. With $\gamma_{0}=0.43$ and $c_{q}=0.8$, the MDLA predictions achieve good agreement not only with PYTHIA simulations but also, as will be shown later, with the ATLAS measurements reported in ref. ATLAS:2019rqw . In details, we first determine $\gamma_{0}$ by fitting the PYTHIA results over the full $p_{T}$ range for gluon jets. Keeping this value of $\gamma_{0}$ fixed, we then extract $c_{q}$ by fitting the PYTHIA results over the full $p_{T}$ range for quark jets. The fitted $c_{q}$ turns out larger than the chosen value of $2/3$ in ref. Dokshitzer:2025fky , reflecting quantitative differences between the two results.
@@
Note that setting the momentum scale in $\alpha_{s}$ to $Q=Rp_{T}$, using LHAPDF Buckley:2014ana , we obtain $\gamma_{0}$ values in the range 0.41–0.51 across the plotted $p_{T}$ interval. [^2] In refs. Dokshitzer:2025owq ; Dokshitzer:2025fky , the two-loop anomalous dimension $\gamma(\alpha_{s})$ is employed in place of $\gamma_{0}$. It is defined as $\gamma(\alpha_{s})=\gamma_{0}-\left[{\beta_{0}}/{4}+{10n_{f}}/{(3N_{c}^{2})}\right]{\alpha_{s}}/{(2\pi)}$, which yields values in the range 0.36–0.43 for $n_{f}=5$. Our choice of $\gamma_{0}=0.43$ indeed lies in this range although allowing for this running produces a broader spread than is seen in the PYTHIA results. Since the MDLA scaling functions with $\gamma_{0}=0.43$ and $c_{q}=0.8$ can very well describe the experimental data in ref. ATLAS:2019rqw , we proceed with this phenomenological approach and defer a more refined analysis beyond DLA to future work.
@@
Similar to the DLA results (see, e.g., ref. Bassetto:1987fq ), we find that the MDLA KNO functions are well approximated by negative binomial distributions (NBDs) of the form
@@
where $k$ serves as a fit parameter (see also ref. Dokshitzer:1993dc ). The right panel of figure 2 demonstrates this agreement, where the MDLA results are reproduced using $k=5.4$ for quark jets and $k=8.5$ for gluon jets, exhibiting only minor deviations.
@@ 3 Inclusive multiplicity distributions in $pp$ collisions at the LHC
In $pp$ collisions, distinguishing quark and gluon jets to study their multiplicity distributions is a nontrivial challenge Marzani:2019hun ; Larkoski:2024uoc , which we defer to the next section. As an alternative, the validity of KNO scaling can be inferred through inclusive multiplicity measurements of QCD jets Germano:2024ier ; Duan:2025ngi ; Dokshitzer:2025owq . In this section, we compare our theoretical calculations, based on the MDLA scaling functions evaluated above, with ATLAS measurements of the inclusive multiplicity distributions of the two leading jets at $\sqrt{s}=13$ TeV ATLAS:2019rqw .
@@ 3.1 Theoretical calculations
Inclusive multiplicity distributions inside jets in $pp$ collisions have been measured at the LHC ATLAS:2011eid ; ATLAS:2019rqw . These inclusive distributions can be represented as a combination of contributions from quark and gluon jets Duan:2025ngi ; Dokshitzer:2025owq :
@@
[…]
@@
[…]
@@
[…]

### v2

@@ 2.2 KNO scaling within QCD jets
In the asymptotic limit $Q\to\infty$, KNO scaling in multiplicity distributions within QCD jets naturally follows the definition of the GFs Konishi et al. (1979); Bassetto et al. (1980); Dokshitzer et al. (1983, 1991); Dremin and Gary (2001). As previously verified in ref. Duan et al. (2025), $\bar{n}_{a}(Q)P_{a}(n,Q)$ for partons within DLA, given in eq. (2.1.2), indeed converge to universal scaling functions for quark and gluon jets respectively for $Q\gg Q_{0}$. However, they are quite different from those for hadrons in PYTHIA simulations and the LHC data. In this subsection, we evaluate the KNO scaling functions within MDLA for both quark and gluon jets, improving the DLA results Dokshitzer et al. (1983); Bassetto (1988); Dokshitzer et al. (1991) by imposing energy conservation Dokshitzer (1993); Dokshitzer and Webber (2025a, b).
@@ 2.2.1 Definition of KNO scaling functions for quark and gluon jets
Following refs. Dokshitzer et al. (1983, 1991), we first outline the definition of, and the method for evaluating, the KNO scaling functions for QCD jets, which holds exactly in the asymptotic limit $Q\to\infty$. Setting $u=\exp\{-\beta/\bar{n}_{a}\}$ in the definition of GFs in eq. (2) and taking the limit $Q\to\infty$ in $Z_{a}$ yields
@@
$\equiv\int_{0}^{\infty}dx~\Psi_{a}(x)e^{-\beta x}=\sum\limits_{k=0}^{\infty}\frac{(-\beta)^{k}}{k!}f_{a}^{(k)}$ (21)
@@ 2.2.2 KNO scaling functions within MDLA
As we are interested in the large $Q$ limit, $n^{(m)}_{a}(Q)$ can be expressed in terms of $f_{a}^{(m)}$ and $\bar{n}_{a}(Q)$ according to eq. (28) in the above equation. Following ref. Dokshitzer (1993), we consider only the corrections arising from energy conservation, while neglecting the effects of the running coupling and higher-order corrections to the anomalous dimension $\gamma_{0}$. In this case, one can approximate the mean multiplicities as
@@
where the digamma function $\psi(z)\equiv\Gamma^{\prime}(z)/\Gamma(z)$. Following ref. Dokshitzer (1993), we neglect the $\psi$ terms. Then, starting with $f_{a}^{(0)}=f_{a}^{(1)}=1$, one can iteratively obtain, for $m\geq 2$,
$f^{(m)}_{g}$ $=\frac{\gamma_{0}m}{m^{2}-1}\sum\limits_{k=1}^{m-1}C_{m}^{k}\frac{\Gamma(\gamma_{0}k)\Gamma(\gamma_{0}(m-k)+1)}{\Gamma(\gamma_{0}m+1)}f^{(k)}_{g}f^{(m-k)}_{g}$,
@@
The DLA results Dokshitzer et al. (1983, 1991) are recovered in the limit $\gamma_{0}\to 0$.
Given the expressions $f_{a}^{(m)}$ above, the KNO scaling functions $\Psi_{a}(x)$ can be presented using an expansion in Laguerre polynomials Bassetto (1988), which obey the relations
@@
Figure 1: KNO scaling functions for quark and gluon jets. Left: Verification of KNO scaling within DLA by directly evaluating $P_{a}(n)$ according to eq. (2.1.2) with $Q_{0}=0.5$ GeV, using a fixed coupling $\alpha_{s}=0.2$, over a $p_{T}$ range from 0.1 to 2.5 TeV (solid), and comparing them to the DLA KNO scaling functions derived in the asymptotic limit $Q\to\infty$ (dashed). Within this $p_{T}$ range, we observe that the curves of $\bar{n}_{a}P_{a}(n)$ for different $p_{T}$ still show some dependence on the value of $\alpha_{s}$. With the chosen value, they all converge to the DLA KNO scaling functions. Right: Comparison of the MDLA KNO scaling functions evaluated from eq. (30) (solid) with the QCD-inspired forms in ref. Dokshitzer and Webber (2025b) (dashed) for $\gamma_{0}=0.4=\gamma$ and $c_{q}=2/3$.
@@
The left panel of figure 1 verifies the validity of KNO scaling in DLA with a fixed $\alpha_{s}$, which is assumed in the preceding MDLA derivation. This is verified by directly evaluating $P_{a}(n)$, as given in eq. (2.1.2) with $Q_{0}=0.5$ GeV, over a broad $p_{T}$ range from 0.1 to 2.5 TeV. With running coupling, the resulting curves of $\bar{n}_{a}P_{a}(n)$ all converge to some universal functions for quark and gluon jets, respectively, though with slight deviations from the asymptotic DLA scaling functions Duan et al. (2025). With fixed coupling, we find that $\bar{n}_{a}P_{a}(n)$ shows some dependence on the value of $\alpha_{s}$ and the results with $\alpha_{s}=0.2$ converge to the DLA scaling functions, as shown in this plot. Here, the asymptotic KNO functions within DLA are given by eq. (36) with $f_{a}^{(n)}$ evaluated in the limit $\gamma_{0}\to 0$, which agree with the results of refs. (Bassetto, 1988; Dokshitzer, 1993).
@@
The right panel of figure 1 compares the MDLA scaling functions evaluated above with the QCD-inspired expressions recently proposed in ref. Dokshitzer and Webber (2025b). The comparison is shown for $\gamma_{0}=0.4$ Dokshitzer (1993) (with $\gamma=0.4$ in the results of ref.Dokshitzer and Webber (2025b)) and $c_{q}=2/3$ Dokshitzer and Webber (2025b). For both quark and gluon jets, including energy conservation significantly shifts the peak of the scaling functions toward larger values of $x$. These significant discrepancies highlight the need to go beyond DLA to obtain reliable KNO scaling functions. Our results, obtained using Laguerre polynomial expansions, agree with the expression for gluon jets given in ref. Dokshitzer (1993), except in the $x\ll 1$ region, where our method may yield negative values, like the DLA results in ref. Bassetto (1988). Hence, for the remainder of the paper, we adopt the analytic form from ref. Dokshitzer (1993) for gluon jets. In comparison with recently suggested forms in ref. Dokshitzer and Webber (2025b), our results for gluon jets are slightly different from that at $x\lesssim 1$, starting around the peak region, while the differences for quark jets are more pronounced. We find that this difference can be traced back to the fact that the coefficients $f_{q}^{(m)}$, calculated iteratively above, no longer satisfy the relation within DLA: $\Phi_{q}(\beta)=[\Phi_{g}(\beta/c_{q})]^{c_{q}}$, which is used to obtain the quark-jet KNO functions in refs. Dokshitzer and Webber (2025a, b).
@@ 2.3 Comparison with KNO scaling functions using PYTHIA
As shown by the results above, while the mean parton multiplicity and the parton multiplicity distribution functions exhibit a logarithmic dependence on $Q_{0}$, the asymptotic KNO scaling functions remain infrared safe. Consequently, one may not expect logarithmically enhanced corrections when transitioning from the parton level to the hadron level. In this subsection, we compare the above results at the parton level with those for charged particles extracted in PYTHIA simulations Bierlich and others (2022), which exhibit KNO scaling to a very reasonable level in both quark and gluon jets Duan et al. (2025).
@@
In our PYTHIA simulations of $pp$ collisions Duan et al. (2025), the event selection procedure follows the methodology in ref. Aad and others (2019). Particles satisfying $p_{T}>0.5$ GeV and $|\eta|<2.5$ are used to reconstruct the two leading jets via the anti-$k_{t}$ algorithm Cacciari et al. (2008) with a radius parameter $R=0.4$, implemented in the FastJet package Cacciari et al. (2012). The jets are required to satisfy $|\eta|<2.1$, and the leading and subleading jets must have a transverse momentum ratio $p_{T}^{\text{lead}}/p_{T}^{\text{sublead}}<1.5$. The PYTHIA samples are generated using the A14 tune, which has been shown to provide a good description for charged-particle multiplicity distribution observed by ATLAS Aad and others (2019). In our analysis, the charged particles from jet constituents are used to define the multiplicity distributions, while jet $p_{T}$ is reconstructed from all particles inside the jet. In addition, quark and gluon jets are identified based on their spatial proximity to the initiating parton from the underlying $2\rightarrow 2$ hard scattering event.
@@
Figure 2: KNO scaling functions for quark and gluon jets in PYTHIA and MDLA. Left: Charged-particle multiplicity distributions for quark and gluon jets in PYTHIA Duan et al. (2025) (solid) versus MDLA results (dashed) with $\gamma_{0}=0.43$ and $c_{q}=0.8$. Right: Negative binomial distributions (NBD, solid), with $k=5.4$ for quark jets and $k=8.5$ for gluon jets (see eq. (38)) compared to the same MDLA results (dashed).
@@
The left panel of figure 2 compares our MDLA results with PYTHIA simulations that include both MPI and hadronization, facilitating comparison with ATLAS data in the next sections. We find that incorporating energy conservation indeed improves significantly the agreement between the KNO scaling functions and the PYTHIA results for both quark and gluon jets. Inspired by refs. Dokshitzer and Webber (2025a, b), we take $c_{q}$ as a fitting parameter rather than fixing it to the DLA value $C_{F}/C_{A}$. With $\gamma_{0}=0.43$ and $c_{q}=0.8$, the MDLA predictions achieve good agreement not only with PYTHIA simulations but also, as will be shown later, with the ATLAS measurements reported in ref. Aad and others (2019). In details, we first determine $\gamma_{0}$ by fitting the PYTHIA results over the entire $p_{T}$ range for gluon jets. Keeping this value of $\gamma_{0}$ fixed, we then extract $c_{q}$ by fitting the PYTHIA results over the entire $p_{T}$ range for quark jets. The fitted $c_{q}$ turns out larger than the chosen value of $2/3$ in ref. Dokshitzer and Webber (2025b), reflecting quantitative differences between the two results.
@@
Note that setting the momentum scale in $\alpha_{s}$ to $Q=Rp_{T}$, using LHAPDF Buckley et al. (2015), we obtain $\gamma_{0}$ values in the range 0.41–0.51 across the plotted $p_{T}$ interval. [^2] In refs. Dokshitzer and Webber (2025a, b), the two-loop anomalous dimension $\gamma(\alpha_{s})$ is employed in place of $\gamma_{0}$. It is defined as $\gamma(\alpha_{s})=\gamma_{0}-\left[{\beta_{0}}/{4}+{10n_{f}}/{(3N_{c}^{2})}\right]{\alpha_{s}}/{(2\pi)}$, which yields values in the range 0.36–0.43 for $n_{f}=5$. Our choice of $\gamma_{0}=0.43$ indeed lies in this range although allowing for this running produces a broader spread than is seen in the PYTHIA results. Since the MDLA scaling functions with $\gamma_{0}=0.43$ and $c_{q}=0.8$ can very well describe the experimental data in ref. Aad and others (2019), we proceed with this phenomenological approach and defer a more refined analysis beyond DLA to future work.
@@
Similar to the DLA results (see, e.g., ref. Bassetto (1988)), we find that the MDLA KNO functions are well approximated by negative binomial distributions (NBDs) of the form
@@
where $k$ serves as a fit parameter (see also ref. Dokshitzer (1993)). The right panel of figure 2 demonstrates this agreement, where the MDLA results are reproduced using $k=5.4$ for quark jets and $k=8.5$ for gluon jets, exhibiting only minor deviations.
@@ 3 Inclusive multiplicity distributions in $pp$ collisions at the LHC
In $pp$ collisions, distinguishing quark and gluon jets to study their multiplicity distributions is a nontrivial challenge Marzani et al. (2019); Larkoski (2024), which we defer to the next section. As an alternative, the validity of KNO scaling can be inferred through inclusive multiplicity measurements of QCD jets Germano et al. (2024); Duan et al. (2025); Dokshitzer and Webber (2025a). In this section, we compare our theoretical calculations, based on the MDLA scaling functions evaluated above, with ATLAS measurements of the inclusive multiplicity distributions of the two leading jets at $\sqrt{s}=13$ TeV Aad and others (2019).
@@ 3.1 Theoretical calculations
Inclusive multiplicity distributions inside jets in $pp$ collisions have been measured at the LHC Aad and others (2011, 2019). These inclusive distributions can be represented as a combination of contributions from quark and gluon jets Duan et al. (2025); Dokshitzer and Webber (2025a):
@@
$r_{a}\equiv\frac{d\sigma_{a}/dp_{T}}{d\sigma_{q}/dp_{T}+d\sigma_{g}/dp_{T}}$ (40)
denotes the fraction of jets initiated by a parton of type $a$ at a given jet transverse momentum $p_{T}$, and $d\sigma_{a}/dp_{T}$ represents the corresponding differential cross section.
@@
with $x_{T}\equiv 2p_{T}/\sqrt{s}$. The PDFs $f_{a/p}(x_{a},\mu^{2})$ and $f_{b/p}(x_{b},\mu^{2})$ are evaluated at the factorization scale $\mu^{2}=p_{T}^{2}$. The partonic cross sections, denoted as $d\hat{\sigma}_{ab\rightarrow cd}/d\hat{t}$ and found in the standard textbooks such as ref. Peskin and Schroeder (1995), are expressed in terms of the Mandelstam variables:
@@
In our analysis, we explicitly separate the contributions from quark and gluon jet channels by examining each partonic subprocess $ab\rightarrow cd$ and evaluating their corresponding partonic cross sections. This enables us to compute the differential cross sections $d\sigma_{q}/dp_{T}$ for quark jets and $d\sigma_{g}/dp_{T}$ for gluon jets, which in turn allows us to extract the quark and gluon fractions $r_{q}$ and $r_{g}$, as defined in eq. (40). For multiplicity distributions, we utilize the MDLA results as presented in the previous section. Based on the theoretical framework outlined above, we investigate the mean multiplicities and multiplicity probability distributions within jets across a broad jet transverse momentum range of $p_{T}=0.1$$2.5$ TeV in $pp$ collisions at $\sqrt{s}=13$ TeV. Our theoretical predictions are to be compared with results from PYTHIA simulations and experimental measurements from ATLAS Aad and others (2019).